# Global parameterizations of $\pi\pi$ scattering with dispersive constraints: beyond the S0 wave

### v1

@@ III.2.4 The three D0-wave Global Fits
Here, it is worth recalling our pending comment that the inelasticity of Solutions II and III was set to begin at the $K\bar{K}$ threshold. Actually, they were obtained using a D0 wave parameterized as a two-channel K-matrix only above the $K\bar{K}$ threshold. In Fig. 4 we see that the data of Solutions II and III below 1.6 GeV do not have error bars because in [26] uncertainties were only provided above that energy. Consequently, to obtain reasonable error bands, we have fit these data assuming an uncertainty similar to that of Solution I. In addition, it is worth noticing that the data from Solutions II and III overlap below 1.4 GeV, making only one set visible in Fig. 4. Finally, in these two solutions, the elasticity was set to $\eta_{2}^{(0)}(s)=1$ below $K\bar{K}$ threshold. As explained in [26] “in fact, all inelastic effects” are parameterized in the $K\bar{K}$ channel of their two-channel K-matrix.
@@
Nevertheless, we emphasize that, within our formalism, we cannot specify the origin of the inelasticity. We just set a phenomenological threshold based on where the data suggests the process is no longer elastic. For instance, our inelasticity may well be due in part to the $4\pi$ channel, known to play a significant role in the $f_{0}(1270)$ decays. Most likely such a contribution dominates the inelasticity below the $K\bar{K}$ threshold. However, how much of the inelasticity arises from a particular channel cannot be discerned within our approach. We only deal with the total $\eta_{\ell}^{(I)}$. On the positive side, our formalism avoids the model-dependence required to identify how many channels are open and what they are made of.
@@
It is also worth noticing that the phase shift around the $f_{0}(1270)$ of the three Global Fits has a rather similar shape to a BW resonance, but that is not the case for the elasticity, since it is very asymmetric. For illustration, we have also represented in Fig. 4 a BW shape as a green line using the present $f_{0}(1270)$ RPP parameters [86]. Recall that, as explained right after Eq. (25), the branching ratio is $21-\epsilon/$ 2 when naively using a BW. Thus, from Table 2, we see that the $f_{2}(1270)$ branching ratio to two pions is smaller for Global Fits II and III than for Fit I: respectively $0.842\pm 0.013$, $0.839\pm 0.012$ and $0.871\pm 0.008$.
@@
III.3 F-wave parameterization
This wave is attractive and, above 1.2 GeV, is dominated by the $\rho_{3}(1690)$, which has a width of $161\pm 10$ MeV [86]. However, its shape differs significantly from the P and D0 waves, since this is a highly inelastic wave in the resonant region, with a branching ratio to $\pi\pi$ of only $\sim 23\%$. As a result, its phase shift is small up to 1.9 GeV, where data exist. The RPP also lists a second F-wave resonance at 1.99 GeV, but this state is currently omitted from the summary tables. Given this situation— and considering that we will fit scattering data up to 1.9 GeV and use this parameterization inside dispersive integrals only below 1.6 GeV—we do not find it necessary to try to describe this second non-confirmed state.
In [54], the dispersive representation was studied only up to 1.4 GeV, where the role of the F wave was almost insignificant. A very naive, non-resonant parameterization was used to describe the data below 1.4 GeV, along with the scattering length obtained from sum rules. The inelasticity was neglected below 1.4 GeV.
Thus, we propose here a completely new parameterization. It will describe the phase shift data, starting somewhat above 0.9 GeV, as well as the elasticity data that starts around 1.3 GeV. We will also fit the scattering length, $a^{1}_{3}$, and effective range, $b^{1}_{3}$, obtained from the CFD in [54] and listed in Table 10 of Appendix A. In this region, the effect of the resonance at 1.69 GeV is negligible. For this reason, our parameterization will have an elastic low-energy part, described by a conformal expansion, and a high-energy part, an inelastic Breit-Wigner parameterization with a Blatt-Weiskopf barrier, as used in [20].
@@ III.3 F-wave parameterization
where $w_{m}=w(s_{m})$, $w^{\prime}_{m}=w^{\prime}(s_{m})$ and the phase shift, elasticity, and their first derivatives at $s_{m}$ in the previous expressions must be calculated using Eqs. (34) and Eq. (35) below, which provide the expression for the high-energy region, to be explained next. Note that, unlike the waves discussed so far, in this case, we fix the parameters of the low-energy part using values from the high-energy parameterization (provided below). We have adopted this matching scheme because the F wave is much better known in the high-energy region. Consequently, the uncertainty of the threshold parameters from the Global Fit will be larger than from their input (see Table 10).
The region $1.2\,{\mathrm{\,Ge\kern-1.00006ptV}}\leq s^{1/2}\leq 1.9\,{\mathrm{\,Ge\kern-1.00006ptV}}$ lies within three widths of the $\rho_{3}$ resonance. Thus, our parameterization in this range employs an inelastic Breit-Wigner shape with a potential barrier, similar to the approach used by the experimentalists in [20]:
@@
are the familiar Blatt-Weiskopf [87] angular momentum barrier factors. For brevity, we will refer to this functional form as BWBW.
Why not use this simple BWBW shape in the whole range? There are several reasons. First, because it yields threshold parameters that are very inconsistent with those obtained from sum rules with the dispersive representation in [54] or with ChPT. Although it is possible to adjust the numerator of the BW shape—e.g., by promoting the $x_{\rho_{3}}$ parameter to a function of $s$—to reproduce these threshold parameters with a shape similar to our Global Fits, the resulting expression would fail to remain elastic in the elastic region. In particular, elastic unitarity implies $\mathrm{Im}\,t^{(1)}_{3}/\sigma|\,t^{(1)}_{3}|^{2}=1$, and with the naively modified inelastic Breit-Wigner this ratio is off by an order of magnitude, even though $\eta^{(1)}_{3}$ may still be very close to one. This has enormous consequences for the FDRs.
For this reason, our global parameterization has a strictly elastic piece in the elastic region matched to the Breit-Wigner shape in the resonance region. These observations will also drive our choice of parameterization for the G0 wave.
Figure 5: F-wave phase shift (top) and elasticity (bottom). We show our unconstrained and dispersively constrained Global Fit I, together with the CFD from [54]. Our new parameterization fits the Hyams et al. data from [20]. The data from Protopopescu et al. [24] are only included for completeness. The red vertical line stands at the allowed onset of the inelasticity, whereas the black one marks the energy of the matching with the BWBW parameterization Eqs. (34) and (35). For illustration, we also show a BWBW curve with the RPP [86] estimates for the mass and width.
@@ III.3.1 F-wave Global Fit and parameters
This parameterization is entirely new. Therefore, we do not use CFD as input at low energies. Instead, we use the threshold parameter values $a^{(1)}_{3}$ and $b^{(1)}_{3}$ provided in Appendix A. In addition, we fit the CERN-Munich scattering data of Hyams et al. [20], shown with solid symbols in Fig. 5. This wave is very small, and its phase shift remains below $2^{\circ}$ up to 1.1 GeV, since the influence of the $\rho_{3}(1690)$ starts to be felt. Our dispersively constrained Global Fit I is shown in the same figure and its parameters are listed in Table 3. We also show the unconstrained Global Fit I. Although it looks qualitatively very similar to the constrained one, it is quite apart in terms of deviations.
Since we fit the scattering data, our $\rho_{3}(1690)$ resonance mass and width are compatible with those obtained from the phase shift analysis by the CERN-Munich collaboration [20], i.e., $m_{\rho_{3}}=1713\pm 4\,$MeV, $\Gamma_{\rho_{3}}=228\pm 10\,$MeV, and $x_{\rho_{3}}=0.26\pm 0.02$. Nevertheless, our Blatt-Weisskopf phenomenological radius is smaller than their value: $R_{\rho_{3}}=6.38\pm 0.44\,{\mathrm{\,Ge\kern-1.00006ptV}}^{-1}$. Overall, our uncertainties are smaller because we impose consistency with the elastic region and the threshold parameters we use as input are very precise.
Note, however, that the mass and width differ by several MeV from the RPP values, dominated by the CERN-Munich BWBW analysis of their $\pi N\to\pi\pi N^{\prime}$ momentum distributions. To illustrate this difference, Fig. 5 shows a naive BWBW shape using the mass and width values estimated by the RPP. The resulting shape is qualitatively similar to our Global Fit, but the phase shift is slightly displaced and the data are described worse. The width is also smaller and misses all the elasticity data above 1.65 GeV. Since we are analyzing scattering data, we adopt our parameterization and do not fit the RPP estimates.
@@ III.3.2 The three F-wave Global Fits
Figure 6: Comparison among Global Fits I, II, and III for the F wave. Qualitatively, they all display a clear resonance shape, although their parameters are incompatible within their uncertainties. Note that the dispersively constrained Global Fit II lies far away from its corresponding data. Above 1.6 GeV (shaded region) the Global Fits are pure data fits since dispersive constraints are only imposed up to that energy. Note that we only plot our Global Fits up to the round energy value closest to the last data point included in the fit. The data come from [20] (Hyams et al. 73) and [26] (Hyams et al. 75) .
III.4 G0-wave parameterization
@@
Unfortunately, no scattering data are available for this wave. Previous dispersive analyses, which focused on the dispersive representation up to 1.4 GeV, relied on a very simple estimate of its imaginary part [73]. Still, based on our previous experience with the F wave, we expect to achieve a fair description of the $f_{4}(2050)$ resonance by describing it with a Breit-Wigner-like form, using its RPP mass and width as input. This will be then matched at low energies to a conformal expansion, where we will impose the scattering length obtained from sum rules using previous dispersive analyses [73].
@@
At intermediate energies, $1.4\,{\mathrm{\,Ge\kern-1.00006ptV}}\leq s^{1/2}\leq 2.25{\mathrm{\,Ge\kern-1.00006ptV}}$, we use a Breit-Wigner with a potential barrier for the $f_{4}(2050)$ resonance:
@@
$B_{1}$ $=\frac{1}{m_{\pi}^{8}w^{\prime}_{m}}\left(\sigma(s)k(s)^{8}\cot\delta_{4}^{(0)}(s)\right)\bigg{|}^{\prime}_{s=s_{m}},$
@@ III.4.1 G0-wave Global Fit and parameters
As commented before, there are no scattering data for this wave. Thus our input consists of the $a^{(0)}_{4}$ scattering length provided in Table 10 in appendix A, along with the mass and width of the $f_{4}(2050)$ resonance taken from the weighted average of their determinations listed in the RPP [86] with two pions in the final state. This yields $m_{f_{4}}=2030\pm 23{\mathrm{\,Me\kern-1.00006ptV}}$ and $\Gamma_{f_{4}}=248\pm 80{\mathrm{\,Me\kern-1.00006ptV}}$. The central values are consistent within uncertainties with the central RPP averages but have uncertainties roughly two to four times larger, which we consider a very conservative estimate. The input value for $x_{4}=0.170\pm 0.030$ was also taken from the RPP but conservatively doubling again its estimated uncertainty. Finally, the value of $R_{f_{4}}$ does not affect the resonant shape much, and given the simplicity of the parameterization, it gets completely fixed from the high precision of the scattering length and the differentiable matching conditions.
@@
In Fig. 7, we show the resulting phase shift (top panel) and elasticity (bottom panel) from the Global Fit I. As already commented, there are no scattering data available for comparison. For illustration, we show in the figure a naive Breit-Wigner shape with the mass and width averages of the RPP. The difference in the central curve arises because we consider only width determinations from two-pion final states, we include Blatt-Weiskopf factors and impose a matching to the scattering length. Still, as far as one width away from its peak, this simple shape is pretty consistent with our Global Fit within uncertainties.
@@
Figure 7: G0-wave phase shift (top) and elasticity (bottom). The red vertical line stands at the onset of the inelasticity, while the black one marks the energy of the matching between the low-energy and the BWBW parameterizations, Eqs. (38) and (39). The unconstrained fit is not visible because it is identical to the constrained one. For illustration, we show a naive Breit-Wigner with the RPP parameters in the region within one width on either side of its mass. Above 1.6 GeV (shaded region), there are no dispersive constraints.
@@
The G0-wave parameters, obtained after matching with the low energy parameterizations and using the dispersion relations as constraints, are given in Table 4. They are very compatible with the input values and those in the RPP. The value of $R_{f_{4}}$ falls within the bulk of values used in the literature for different resonances, which is of the order of a few ${\mathrm{\,Ge\kern-1.00006ptV}}^{-1}$. It is indeed very close to the one we found for the F wave and also in very good agreement with the values found for the $\rho(770)$ in the literature, typically $4-5{\mathrm{\,Ge\kern-1.00006ptV}}^{-1}$ [88, 20, 28, 89, 90]. However, given that with this parameter we are basically fitting just one high-precision datum, which is the scattering length, The resulting uncertainty is smaller than 0.5 MeV, and, in practice, $R_{f_{4}}$ becomes a fixed constant in our phenomenological parameterization. We have nevertheless allowed it to vary when imposing the dispersive constraints.

### v2

@@ III.2.4 The three D0-wave Global Fits
Here, it is worth recalling our pending comment that the inelasticity of Solutions II and III was set to begin at the $K\bar{K}$ threshold. Actually, they were obtained using a D0 wave parametrized as a two-channel K-matrix only above the $K\bar{K}$ threshold. In Fig. 4 we see that the data of Solutions II and III below 1.6 GeV do not have error bars because in [26] uncertainties were only provided above that energy. Consequently, to obtain reasonable error bands, we have fit these data assuming an uncertainty similar to that of Solution I. In addition, it is worth noticing that the data from Solutions II and III overlap below 1.4 GeV, making only one set visible in Fig. 4. Finally, in these two solutions, the elasticity was set to $\eta_{2}^{(0)}(s)=1$ below $K\bar{K}$ threshold. As explained in [26] “in fact, all inelastic effects” are parametrized in the $K\bar{K}$ channel of their two-channel K-matrix.
@@
Nevertheless, we emphasize that, within our formalism, we cannot specify the origin of the inelasticity. We just set a phenomenological threshold based on where the data suggests the process is no longer elastic. For instance, our inelasticity may well be due in part to the $4\pi$ channel, known to play a significant role in the $f_{2}(1270)$ decays. Most likely such a contribution dominates the inelasticity below the $K\bar{K}$ threshold. However, how much of the inelasticity arises from a particular channel cannot be discerned within our approach. We only deal with the total $\eta_{\ell}^{(I)}$. On the positive side, our formalism avoids the model-dependence required to identify how many channels are open and what they are made of.
@@
It is also worth noticing that the phase shift around the $f_{2}(1270)$ of the three Global Fits has a rather similar shape to a BW resonance, but that is not the case for the elasticity, since it is very asymmetric. For illustration, we have also represented in Fig. 4 a BW shape as a green line using the present $f_{2}(1270)$ RPP parameters [90]. Recall that, as explained right after Eq. (25), the branching ratio is $21-\epsilon/$ 2 when naively using a BW. Thus, from Table 2, we see that the $f_{2}(1270)$ branching ratio to two pions is smaller for Global Fits II and III than for Fit I: respectively $0.842\pm 0.013$, $0.839\pm 0.012$ and $0.871\pm 0.008$.
@@
III.3 F-wave parametrization
This wave is attractive and, above 1.2 GeV, is dominated by the $\rho_{3}(1690)$, which has a width of $161\pm 10$ MeV [90]. However, its shape differs significantly from the P and D0 waves, since this is a highly inelastic wave in the resonant region, with a branching ratio to $\pi\pi$ of only $\sim 23\%$. As a result, its phase shift is small up to 1.9 GeV, where data exist. The RPP also lists a second F-wave resonance at 1.99 GeV, but this state is currently omitted from the summary tables. Given this situation— and considering that we will fit scattering data up to 1.9 GeV and use this parametrization inside dispersive integrals only below 1.6 GeV—we do not find it necessary to try to describe this second non-confirmed state.
In [54], the dispersive representation was studied only up to 1.4 GeV, where the role of the F wave was almost insignificant. A very naive, non-resonant parametrization was used to describe the data below 1.4 GeV, along with the scattering length obtained from sum rules. The inelasticity was neglected below 1.4 GeV.
Thus, we propose here a completely new parametrization. It will describe the phase shift data, starting somewhat above 0.9 GeV, as well as the elasticity data that starts around 1.3 GeV. We will also fit the scattering length, $a^{(1)}_{3}$, and effective range, $b^{(1)}_{3}$, obtained from the CFD in [54] and listed in Table 10 of Appendix A. In this region, the effect of the resonance at 1.69 GeV is negligible. For this reason, our parametrization will have an elastic low-energy part, described by a conformal expansion, and a high-energy part, an inelastic Breit-Wigner parametrization with a Blatt-Weisskopf barrier, as used in [20].
@@ III.3 F-wave parametrization
where $w_{m}=w(s_{m})$, $w^{\prime}_{m}=w^{\prime}(s_{m})$ and the phase shift, elasticity, and their first derivatives at $s_{m}$ in the previous expressions must be calculated using Eqs. (34) and Eq. (35) below, which provide the expression for the high-energy region, to be explained next. Note that, unlike the waves discussed so far, in this case, we fix the parameters of the low-energy part using values from the high-energy parametrization (provided below). We have adopted this matching scheme because the F wave is much better known in the high-energy region. Consequently, the uncertainty of the threshold parameters from the Global Fit will be larger than from their input (see Table 10).
The region $1.2\,{\mathrm{\,Ge\kern-1.00006ptV}}\leq s^{1/2}\leq 1.9\,{\mathrm{\,Ge\kern-1.00006ptV}}$ lies within three widths of the $\rho_{3}(1690)$ resonance. Thus, our parametrization in this range employs an inelastic Breit-Wigner shape with a potential barrier, similar to the approach used by the experimentalists in [20]:
@@
are the familiar Blatt-Weisskopf [91] angular momentum barrier factors. For brevity, we will refer to this functional form as BWBW. This equation only differs significantly from the naive BW formula, i.e., with a constant width $\Gamma_{\rho_{3}}$ in Eq. (34), for energies far from $m_{\rho_{3}}$.
Why not use this simple BWBW shape in the whole range? There are several reasons. First, even though the $k(s)^{7}$ factor provides the expected threshold leading behavior, it does not account for the subleading contributions. In addition, Eq. (34) corresponds to an inelastic partial wave, although at threshold, it should be elastic. Thus, Eq. (34) alone yields threshold parameters that are very inconsistent with those obtained from sum rules with the dispersive representation in [54] or with ChPT. Although it is possible to adjust the numerator of the BW shape—e.g., by promoting the $x_{\rho_{3}}$ parameter to a function of $s$—to reproduce these threshold parameters with a shape similar to our Global Fits, the resulting expression would fail to remain elastic in the elastic region. In particular, elastic unitarity implies $\mathrm{Im}\,t^{(1)}_{3}/\sigma|\,t^{(1)}_{3}|^{2}=1$, and with such a naively modified inelastic BWBW formula, this ratio is off by an order of magnitude, even though $\eta^{(1)}_{3}$ may still be very close to one. This has enormous consequences for the FDRs.
For these reasons, our global parametrization has a strictly elastic piece in the elastic region matched to the BWBW shape in the resonance region. These observations will also drive our choice of parametrization for the G0 wave.
Figure 5: F-wave phase shift (top) and elasticity (bottom). We show our unconstrained and dispersively constrained Global Fit I, together with the CFD from [54]. Our new parametrization fits the Hyams et al. data from [20]. The data from Protopopescu et al. [24] are only included for completeness. The red vertical line stands at the allowed onset of the inelasticity, whereas the black one marks the energy of the matching with the BWBW parametrization Eqs. (34) and (35). For illustration, we also show a BWBW curve with the RPP [90] estimates for the mass and width. We also illustrate the inability of a simple BW parametrization to describe the data and provide a good low-energy description.
@@ III.3.1 F-wave Global Fit and parameters
This parametrization is entirely new. Therefore, we do not use CFD as input at low energies. Instead, we use the threshold parameter values $a^{(1)}_{3}$ and $b^{(1)}_{3}$ provided in Appendix A. In addition, we fit the CERN-Munich scattering data of Hyams et al. [20], shown with solid symbols in Fig. 5. This wave is very small, and its phase shift remains below $2^{\circ}$ up to 1.1 GeV, where the influence of the $\rho_{3}(1690)$ starts to be felt. Our dispersively constrained Global Fit I is shown in the same figure and its parameters are listed in Table 3. We also show the unconstrained Global Fit I. Although it looks qualitatively very similar to the constrained one, it is quite apart in terms of deviations. In addition, we show an unconstrained BW fit to illustrate the need for a more elaborate description, like the BWBW form. The naive BW formula roughly reproduces the minimum in the elasticity and the zero in the phase shift. However, it fails to describe the shape of the phase shift, the width of the elasticity dip—which comes out too narrow—and it also has a disastrous low-energy behavior, impossible to match smoothly to the conformal parametrization.
Since we fit the scattering data, our $\rho_{3}(1690)$ resonance mass and width are compatible with those obtained from the phase shift analysis by the CERN-Munich collaboration [20], i.e., $m_{\rho_{3}}=1713\pm 4\,$MeV, $\Gamma_{\rho_{3}}=228\pm 10\,$MeV, and $x_{\rho_{3}}=0.26\pm 0.02$. Nevertheless, our Blatt-Weisskopf phenomenological radius is smaller than their value: $R_{\rho_{3}}=6.38\pm 0.44\,{\mathrm{\,Ge\kern-1.00006ptV}}^{-1}$. Overall, our uncertainties are smaller because we impose consistency with the elastic region and the threshold parameters we use as input are very precise. The most extreme case is the uncertainty of $R_{\rho_{3}}$, which comes out about twenty times smaller than the CERN-Munich result. This could be expected since, as explained above, well below the resonance nominal mass, this parameter controls the BWBW behavior and dominates its uncertainty. A $R_{\rho_{3}}$ variation from its central value as large as that of the CERN-Munich one standard deviation would result in a scattering length more than ten deviations away from the input value.
Note, however, that the mass and width differ by several MeV from the RPP values, dominated by the CERN-Munich BWBW analysis of their $\pi N\to\pi\pi N^{\prime}$ momentum distributions. To illustrate this difference, Fig. 5 shows a naive BWBW shape using the mass and width values estimated by the RPP. The resulting shape is qualitatively similar to our Global Fit, but the phase shift is slightly displaced, and the data are described worse. The width is also smaller and misses all the elasticity data above 1.65 GeV. Since we are analyzing scattering data, we adopt our parametrization and do not fit the RPP estimates.
@@ III.3.2 The three F-wave Global Fits
Figure 6: Comparison among Global Fits I, II, and III for the F wave. Qualitatively, they all display a clear resonance shape, although their parameters are incompatible within their uncertainties. Note that the dispersively constrained Global Fit II lies far away from its corresponding data. Above 1.6 GeV (shaded region), the Global Fits are pure data fits since dispersive constraints are only imposed up to that energy. Note that we only plot our Global Fits up to the round energy value closest to the last data point included in the fit. The data come from [20] (Hyams et al. 73) and [26] (Hyams et al. 75).
III.4 G0-wave parametrization
@@
Unfortunately, no scattering data are available for this wave. Previous dispersive analyses, which focused on the dispersive representation up to 1.4 GeV, relied on a very simple estimate of its imaginary part [73]. Still, based on our previous experience with the F wave, we expect to achieve a fair description of the $f_{4}(2050)$ resonance by describing it with a Breit-Wigner-like form, using its RPP mass and width as input. This will then be matched at low energies to a conformal expansion, where we will impose the scattering length obtained from sum rules using previous dispersive analyses [73].
@@
Above 1.4 GeV, we use a Breit-Wigner with a potential barrier for the $f_{4}(2050)$ resonance:
@@
$B_{1}$ $=\frac{1}{m_{\pi}^{8}w^{\prime}_{m}}\left(\sigma(s)k(s)^{8}\cot\delta_{4}^{(0)}(s)\right)^{\prime}\bigg{|}_{s=s_{m}},$
@@ III.4.1 G0-wave Global Fit and parameters
As commented before, there are no scattering data for this wave. Thus our input consists of the $a^{(0)}_{4}$ scattering length provided in Table 10 in Appendix A, along with the mass and width of the $f_{4}(2050)$ resonance taken from the weighted average of their determinations listed in the RPP [90] with two pions in the final state. This yields $m_{f_{4}}=2030\pm 23{\mathrm{\,Me\kern-1.00006ptV}}$ and $\Gamma_{f_{4}}=248\pm 80{\mathrm{\,Me\kern-1.00006ptV}}$. The central values are consistent within uncertainties with the central RPP averages but have uncertainties roughly two to four times larger, which we consider a very conservative estimate. The input value for $x_{f_{4}}=0.17\pm 0.03$ was also taken from the RPP but conservatively doubling again its estimated uncertainty. Finally, the value of $R_{f_{4}}$ does not affect the resonant shape much, and given the simplicity of the parametrization, it gets completely fixed from the high precision of the scattering length and the differentiable matching conditions.
@@
In Fig. 7, we show the resulting phase shift (top panel) and elasticity (bottom panel) from the Global Fit I. As already commented, there are no scattering data available for comparison. For illustration, we show in the figure a naive Breit-Wigner shape with the mass and width averages of the RPP. The difference in the central curve arises because we consider only width determinations from two-pion final states, we include Blatt-Weisskopf factors and impose a matching to the scattering length. Still, as far as one width away from its peak, this simple shape is pretty consistent with our Global Fit within uncertainties.
@@
Figure 7: G0-wave phase shift (top) and elasticity (bottom). The red vertical line stands at the onset of the inelasticity, while the black one marks the energy of the matching between the low-energy and the BWBW parametrizations, Eqs. (38) and (39). The unconstrained fit is not visible because it is identical to the constrained one. For illustration, we show a naive Breit-Wigner with the RPP parameters in the region within one width on either side of its mass. Above 1.6 GeV (shaded region), there are no dispersive constraints.
@@
The G0-wave parameters, obtained after matching with the low energy parametrizations and using the dispersion relations as constraints, are given in Table 4. They are very compatible with the input values and those in the RPP. The value of $R_{f_{4}}$ falls within the bulk of values used in the literature for different resonances, which is of the order of a few ${\mathrm{\,Ge\kern-1.00006ptV}}^{-1}$. It is indeed very close to the one we found for the F wave and also in very good agreement with the values found for the $\rho(770)$ in the literature, typically $4-5{\mathrm{\,Ge\kern-1.00006ptV}}^{-1}$ [92, 20, 28, 93, 94]. However, given that with this parameter we are basically fitting just one high-precision datum, which is the scattering length, The resulting uncertainty is smaller than 0.5 MeV, and, in practice, $R_{f_{4}}$ becomes a fixed constant in our phenomenological parametrization. We have nevertheless allowed it to vary when imposing the dispersive constraints.